convert the raw score of 50 to a z-score


1
Expert's answer
2021-05-07T09:13:35-0400

To calculate a z-score from raw score, subtract the mean from the raw score and divide that answer by the standard deviation.


"z=\dfrac{X-\mu}{\sigma}"

where "X" = Raw score

"\mu =" mean

"\sigma=" Standard Deviation

So, Here let "\mu =" 40

and "\sigma =10"


So, Z- score

"z=\dfrac{X-\mu}{\sigma}\\z=\dfrac{50-40}{10}\\\Rightarrow z=1"

Hence, Z- score =1